If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are
I have the answer but cannot figure out how to work problem. Please help

Answer by Alan3354(69443)

If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are
------------
72/3 = 24 mi/hr
60/6 = 10 mi/hr
-----
The boat's speed is the average = (24+10)/2
= 17 mi/hr
---------
The current is the difference = 24-17 = 7 mi/hr
----------
Or, the hard way:
b = boat's speed
c = current
------
b+c = 24
b-c = 10
---------- Add
2b = 34
b = 17 mi/hr
etc.